An elevator is descending down with uniform acceleration. To measure it a person in elevator dropped a coin which is 6 feet above the floor of elevator at the time of elevator starts. The person found the coin touches the ground in 1second.From these data find the acceleration of the elevator. 

Grade:12th pass

1 Answers

Arun
4 years ago
For elevator and coin : Initial velocity=u=0m/s
As the elevator descends  with uniform acceleration "a" the coin has to move more distance than 1.8cm to strike floor.
time taken =t=1 sec
Scoin=Sc=ut+1/2at²
Sc= 0+1/2at²
=1/2 xg x1²
Sc=g/2
Se=ut+1/2 at²
=0+1/2 a(1)²
Se=a/2
Total distance travelled by coin is given by :
1.8 +1/2 a =1/2g
1.8 +a/2 =g/2
1.8+a/2=9.8/2
a/2=4.9 -1.8=3.1m
a=6.2 m/s2
a= 6.2 x 3.28 = 20.34 ft/s²